OpenStudy (anonymous):

giving medal for anwser Find the x-intercept and the y-intercept for the graph of the equation. x-3y=3 a. 3; –2 c. 2; 0 b. 3; –1 d. 4; –1

OpenStudy (anonymous):

well wheres the picture

OpenStudy (anonymous):

theres no pic.

OpenStudy (maddyx):

x= 3 y= (-)1 answer is (3, -1)

OpenStudy (maddyx):

for x-intercept, put y=0 for y-intercept, put x=0 get the co-ordinates

OpenStudy (anonymous):

@maddyx can u help me with like 3 more like this ?

OpenStudy (maddyx):

@leahhbe yes! sure!!

OpenStudy (anonymous):

B 3,-1

OpenStudy (anonymous):

x-3y=3 divide by 3 \[\frac{ x }{3 }-\frac{ 3y }{3 }=1\] \[\frac{ x }{ 3 }+\frac{ y }{ -1 }=1\] x intercept =3,y intercept=-1

OpenStudy (maddyx):

@surjithayer this is the real method

OpenStudy (anonymous):

yw